OpenStudy (anonymous):

Make p the subject of this equation: k=a/2(1/2-p)

OpenStudy (anonymous):

Does that mean solve for p? What does making p the subect mean?

OpenStudy (anonymous):

I would divde both sides to get k/(a/2) = (1/2-p) then 2k/a-1/2 = 1/2-p -1/2 to get p = (2k/a) - (1/2) is this what you are looking for?

OpenStudy (anonymous):

hi mathdood - yes solve for p

OpenStudy (anonymous):

Alright it should be what I said p = (2k/a)-(1/2) you just use algebra do you understand how I got this answer?

OpenStudy (anonymous):

yes - thanks so much mathdood. I am new to studying this and need a pointer in the right direction. This site will be very helpful. Thanks again.